OpenStudy (anonymous):

13 out of 25 what is my grade

OpenStudy (anonymous):

52%

OpenStudy (anonymous):

so an F

OpenStudy (anonymous):

52=F

OpenStudy (texaschic101):

if there is 25 questions, then each question counts as 4 points. You got 13 correct....so thats 13 x 4 = 52....thats an F..sorry

sammixboo (sammixboo):

\(\sf \large \color{red}{\frac{13}{25}}\) To figure out your grade you can make multiply the fraction by \(\sf \color{red}{4}\) to get a denominator of 100 and a numerator of 52. \(\sf \large \color{red}{\frac{13*4}{25*4}}\color{black}{=}\color{red}{\frac{52}{100}}\) And \(\sf \large \color{red}{\frac{52}{100}}\) is \(\sf ~ \color{red}{52}\)%

OpenStudy (anonymous):

thanks

sammixboo (sammixboo):

No problem. were you asking about what your grade will be, or was it a problem?

OpenStudy (anonymous):

wat my grade would be but i got a f